# Scaling Laws for Online Advertisement Retrieval

### v1

@@ 4. Scaling Laws of Budget and Revenue
In this section, we verify whether MLPs exhibit scaling laws in the context of online advertisement retrieval, as described in Section 3. We present a lightweight and effective paradigm for identifying the scaling laws of computation budget and revenue. In sub-section 4.1, we introduce $R/R^{*}$, a carefully designed offline metric that is highly linearly correlated with online revenue. In sub-section 4.2, we demonstrate that MLPs exhibit a scaling law between FLOPs and the offline metric $R/R^{*}$. In sub-section 4.3, we propose two computation cost analysis schemes based on model size parameters. Considering the clear relationship between model size parameters and FLOPs, we can approximately determine the expected computing cost and revenue benefits when the model is deployed online with the above work.
4.1. Effective Surrogate Metric for Online Revenue
@@
Traditional offline metrics (e.g. OPA (Wang et al., 2024), Recall, and NDCG) used for online advertising retrieval often ignore the revenue difference of the ground-truth ads and treat them equally. This naturally creates a gap between offline metrics and online revenue. To mitigate this gap, we propose a novel metric named $R/R^{*}$ based on $D_{train}$. The $R/R^{*}_{(i)}$ for impression $i$ in $D_{train}$ can be formulated as Eq 5:
@@
(5) $\begin{split}R/R^{*}_{(i)}=\frac{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}{\sum(\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}[:m;:]*\textbf{ecpm}_{i}^{T})}\end{split}$
@@
where $\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}$ and $\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}$ denote the hard permutation matrices sorted by $\mathcal{M}^{(i,\cdot)}$ and $\textbf{v}_{i}$ respectively. $\textbf{ecpm}_{i}$ is a row vector that represents the expected revenue of each ad of the $i$-th impression. $R/R^{*}$ is the average of $R/R^{*}_{(i)}$ on $D_{train}$. $m$ is the number of ground-truth ads for exposure.
@@
$R/R^{*}$ aligns better with online revenue mainly because it explicitly considers the commercial value (eCPM) of each ad, directly reflects the goal of maximizing revenue, and reduces the gap between offline evaluation and online performance. If the following two assumptions are met, a linear relationship between $R/R^{*}$ and online revenue is likely to exist:
Assumption 1.
@@
The improvement in $R/R^{*}$ for a single pathway is proportional to the improvement in $R/R^{*}$ of the entire stage ensemble’s output set.
Assumption 2.
The post-stage models generally select only the top-ranked ads from the pre-stage models, and different ads with the same eCPM have the same probability for exposure.
In a mature advertising system, Assumption 2 holds approximately true. Assumption 1 also holds approximately true for pathways with larger weights. Therefore, the $R/R^{*}$ of the LTR pathway and the revenue of the Kuaishou advertising system should be approximately linearly related.
To verify the effectiveness of $R/R^{*}$, we conduct extensive online A/B test experiments. We deploy several different models to empirically study the relationship between online revenue and offline metrics, which differ in feature engineering, surrogate loss, model structure, and model size. Due to the limitations on experimental traffic, we can only run up to 5 A/B test groups simultaneously, which means some experimental groups were launched on different days. The online experiments corresponding to each model use at least 5% of the traffic. We collect each model’s online revenue from the A/B test platform and offline metrics from the training log. Figures 3, 4, 5 and 6 demonstrates the relationship between the online revenue and offline metrics, including our proposed $R/R^{*}$ and traditional retrieval metrics, namely $NDCG$, $Recall$ and $OPA$. To protect commercial secrets, we applied a linear transformation to the original data, and the y-axis represents the relative increase in revenue rather than the absolute value. These operations do not affect the conclusions regarding the linear relationship. Each point in these figures represents data from one day of a model being live. It is evident that:
  • $R/R^{*}$ exhibits the strongest linear correlation with online revenue, with an $R^{2}$ of 0.902, indicating that $R/R^{*}$ is a more advanced offline metric.
  • $NDCG$ exhibits an acceptable linear correlation with online revenue, with an $R^{2}$ of approximately 0.795. Since ads at the same rank in NDCG have the same weight, NDCG is likely a good metric only in systems where ads at the same position have similar eCPM. When training data does not record eCPM information, $NDCG$ would be an alternative to $R/R^{*}$.
  • $Recall$ and $OPA$ exhibit poor linear correlations with online revenue, with $R^{2}$ values of 0.509 and 0.140 respectively, suggesting that they are not good indicators of online revenue.
Figure 3. The relationship between $R/R^{*}$ and the relative increase in online revenue. The $R^{2}$ value is 0.902. The average deviation for the fitted data points is 0.243, and the maximum deviation is 0.483.
Figure 4. The relationship between $NDCG$ and the relative increase in online revenue. The $R^{2}$ value is 0.793. The average deviation for the fitted data points is 0.311, and the maximum deviation is 0.819.
Figure 5. The relationship between $Recall$ and the relative increase in online revenue. The $R^{2}$ value is 0.509. The average deviation for the fitted data points is 0.534, and the maximum deviation is 1.139.
Figure 6. The relationship between $OPA$ and the relative increase in online revenue. The $R^{2}$ value is 0.140. The average deviation for the fitted data points is 0.634, and the maximum deviation is 1.690.
4.2. Scaling Laws of FLOPs and Offline Metric
@@
In this section, we verify whether there is a scaling law for the model between computational effort and online revenue under the setting described in Section 3. Thanks to the strong linear correlation between $R/R^{*}$ and online revenue, we can transform the process of identifying the scaling law from online experiments to offline experiments. We trained 10 models with different MLP sizes offline, each for 10 days to ensure convergence. We then collected their converged $R/R^{*}$ metrics and calculated the FLOPs for each model to verify whether there is a scaling law for MLP models under our setting. We verify whether the collected data follows the Broken Neural Scaling Law (BNSL), which is a statistical law that has been widely validated by Caballero et al.((2023)(Caballero et al., 2023) for many scaling behaviors of artificial neural networks in CV and NLP tasks. BNSL can be formulated as shown in Eq 6:
(6) $\begin{split}R/R^{*}&=a+(b*FLOPs^{-c_{0}})\prod_{i=1}^{t}(1+(\frac{FLOPs}{d_{i}})^{1/f_{i}})^{(-c_{i}*f_{i})}\end{split}$
where $a$,$b$,$c_{0}$,$c_{i}$,$f_{i}$ are parameters and $t$ is hyper-parameter of BNSL. $t$ characterizes the degrees of freedom of BNSL, and in practice, we typically take it as 6. The scaling factor is the FLOPs, and the dependent variable is our proposed offline metric $R/R^{*}$. We use basin-hopping (Wales and Doye, 1997) to fit the BNSL.
@@
Figure 7. Scaling Performance regarding the FLOPs and $R/R^{*}$ of models. The size of the MLP network and the deviation between the fitted curve and the true values (referred to as the ”error”) are annotated in the figure. The $R^{2}$ value of the curve fitting is 0.996. Here [1024, 256, 256, 256, 1] indicates an MLP model with 5 layers, where the output sizes are 1024, 256, 256, 256, and 1, respectively. The total dimensions of input features for all models are 3328.
@@
Figure 7 illustrates the curve obtained after fitting the BNSL, along with the original data points. The x-axis represents the scaling factor (FLOPs), and the y-axis represents the dependent variable ($R/R^{*}$). Each data point corresponds to an MLP model of a different size, with the number of layers and units per layer annotated in the figure. The $R^{2}$ value of the curve fitting is 0.996, indicating a strong correlation between the predicted values and the observed data points. Therefore, it provides strong evidence that there is indeed a Broken Neural Scaling Law for the MLP models under the typical scenario (referred to Section 3) of online advertisement retrieval. With this scaling law, we can predict or understand the scaling behavior of the MLP models.
It is important to note that the FLOPs in Figure 7 refer to the computational cost per user-ad pair (without customized optimization, e.g. the first-layer optimization), and the training data includes about 200 billion samples (namely user-ad pairs) and 20 billion impressions. We set the maximum FLOPs for the model to 143M FLOPs. The FLOPs range is not as wide as in NLP (Kaplan et al., 2020; Hoffmann et al., 2022). This is mainly because 143M FLOPs is about 20 times the FLOPs of our online base model, and a mature commercial system typically cannot easily allocate such a large amount of computing resources in the short term. Additionally, from the scaling laws validated in other scenarios, the marginal returns of increasing computational cost tend to decrease. Therefore, from the perspective of commercial applications of scaling law in advertising systems, which strongly focus on ROI, predicting model performance within the computational range of 0 to 143M FLOPs should be sufficient.

### v2

@@ 3 Scaling Laws of Budget and Revenue
In this section, we present a lightweight paradigm for identifying the scaling laws of computation budget (namely, machine cost) and revenue. We also verify whether MLP [1], DSSM [27], and Transformer [45] exhibit scaling laws in the context of online advertisement retrieval. In sub-section 3.1, we first introduce $R/R^{*}$, a carefully designed offline metric that is highly linearly correlated with online revenue. In sub-section 3.2, we then demonstrate that DSSMs, MLPs, and Transformers exhibit a broken neural scaling law [9] between FLOPs and the offline metric $R/R^{*}$. In sub-section 3.3, we finally propose a simulation algorithm to estimate machine cost based on model size parameters. Leveraging the $R/R^{*}$ metric, the observed scaling laws, and the cost simulation algorithm, we are able to reliably estimate online machine costs and revenue gains from offline model configurations.
3.1 Effective Surrogate Metric for Online Revenue
@@
Traditional offline metrics (e.g., OPA [49], Recall, and NDCG) used for online advertising retrieval often ignore the revenue difference of the ground-truth ads and treat them equally. This naturally creates a gap between offline metrics and online revenue. To mitigate this gap, we propose a novel metric named $R/R^{*}$ based on $D_{train}$. The $R/R^{*}_{(i)}$ for impression $i$ in $D_{train}$ can be formulated as Eq 2:
@@
$\begin{split}R/R_{(i)}^{*}(m)=\frac{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}{\sum(\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}[:m;:]*\textbf{ecpm}_{i}^{T})}\end{split}$ (2)
@@
where $\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}$ and $\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}$ denote the hard permutation matrices [20] sorted by $\mathcal{M}^{(i,\cdot)}$ and $\textbf{v}_{i}$ respectively. $\mathcal{M}^{(i,\cdot)}$ denotes the prediction vector of model $\mathcal{M}$ for the $i$-th impression of $D_{train}$. When the input vector length is $n$, the hard permutation matrix is an $n\times n$ square matrix. The notation $[:m;:]$ denotes the operation of taking the first $m$ rows of the matrix. $\textbf{ecpm}_{i}$ is a row vector that represents the expected revenue of each ad of the $i$-th impression. $R/R^{*}$ is the average of $R/R^{*}_{(i)}$ on $D_{train}$. $m$ is a hyper-parameter of $R/R^{*}$.
@@
$R/R^{*}$ aligns better with online revenue mainly because it explicitly considers the commercial value (eCPM) of each ad, directly reflects the goal of maximizing revenue, and reduces the gap between offline evaluation and online performance. If the following assumptions are met, we can prove that the R/R* and online revenue have a linear relationship. The proof is in the Appendix B.1.
Assumption 3.1.
The training data $D_{train}$ and online data are independent and identically distributed.
Assumption 3.2.
@@
The improvement in $R/R^{*}$ for a single pathway is proportional to the improvement in $R/R^{*}$ of the entire stage ensemble’s output set. It can be formulated as:
$\begin{split}\Delta R/R^{*}_{ensemble}(m)=\alpha\Delta R/R^{*}_{single}(m)\end{split}$ (3)
where $R/R^{*}_{single}$ are $R/R^{*}_{ensemble}$ the $R/R^{*}$ for single pathway and entire stage, respectively. $\alpha$ is a constant.
Assumption 3.3.
Only the top-$m$ ads from the retrieval stages will be selected by the post-stages for exposure; the exposure probability of $ad_{i}^{j}$ is $p_{i}^{j}$. We define $\beta_{i}^{j}=\frac{\text{eCPM}_{i}^{j}}{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}$ as the normalized contribution of the $j$-th ad (on the top-m sets sorted by retrieval stages) of the $i$-th impression. We assume that $\beta_{i}^{j}$ and $p_{i}^{j}$ are invariant across different impressions, depending only on the position $j$.
Assumption 3.4.
The ranking stage predicts CPMs that are consistent with the true per-impression revenue. Moreover, for each impression $i$, the sum of the top-$M$ predicted CPMs (denoted as $\mathcal{P}_{\textbf{v}_{i}}^{\downarrow}[:m;:]*\textbf{ecpm}^{T}$) is statistically independent of the retrieval model’s performance metric $R/R^{*}_{(i)}$.
In a mature advertising system, we think Assumption ˜3.1, 3.3 and 3.4 holds approximately true. Assumption 3.2 also holds approximately true for pathways with dominant weights. Therefore, the $R/R^{*}$ of the Learning-to-rank pathway (with dominant weight) and the revenue of online advertising systems should be approximately linearly related. To verify the effectiveness of $R/R^{*}$, we also conduct extensive online A/B test experiments. We deploy several different models to empirically study the relationship between online revenue and offline metrics, which differ in feature engineering, surrogate loss, model structure, and model size. Table 1 shows the experimental results. It is evident that $R/R^{*}$ exhibits the strongest linear correlation with online revenue, with an $R^{2}$ of 0.902, indicating that $R/R^{*}$ is a more advanced offline metric. Note that $m$ is 2 in the experiment of Table 1. Due to space limitations, more experimental details are shown in Appendix Sections˜B.2 and B.3.
Table 1: Linear correlation analysis between online revenue and four offline ranking metrics across two stages (Pre-ranking and Matching). $R^{2}$ measures the goodness of fit (higher is better), while average deviation (Avg. Dev) and maximum deviation (Max. Dev) reflect the fitting errors (lower is better). Best results are bolded.
Scenario Correlation Measures Offline Metrics
$R/R^{*}$ NDCG Recall OPA
Pre-ranking $R^{2}\uparrow$ 0.902 0.793 0.509 0.140
[…]
Matching $R^{2}\uparrow$ 0.725 0.315 0.504 0.259
[…]
3.2 Scaling Laws of FLOPs and Offline Metric
@@
In this section, we investigate whether there exist scaling laws for different model architectures between computational effort and online revenue under the setting described in Section 2. Thanks to the strong linear correlation between $R/R^{*}$ and online revenue, we can transform the identification of the scaling laws from costly online A/B tests to efficient offline experiments. We train multiple models with different model sizes offline, and then collected their converged $R/R^{*}$ metrics and calculated the FLOPs for each model to examine the existence of scaling laws for mainstream model architectures (i.e., Transformer [45], MLP [1], and DSSM [27]) under two different scenarios (denoted as $Scene_{1}$ and $Scene_{2}$) in our system. Specifically, we examine whether the collected data follows the Broken Neural Scaling Law (BNSL), a statistical framework proposed by Caballero et al.(2023) and empirically validated across numerous deep learning tasks in computer vision and natural language processing. BNSL is formulated as shown in Eq 4:
$\begin{split}R/R^{*}&=a+(b*FLOPs^{-c_{0}})\prod_{i=1}^{t}(1+(\frac{FLOPs}{d_{i}})^{1/f_{i}})^{(-c_{i}*f_{i})}\end{split}$
[…]
@@
Figure 3: Scaling Performance regarding the FLOPs and $R/R^{*}$ of MLP models under $Scene_{1}$. The size of the MLP network and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure. The $R^{2}$ value of the curve fitting is 0.996. [1024, 256, 256, 256, 1] indicates an MLP model with 5 layers, where the output sizes are 1024, 256, 256, 256, and 1, respectively. The total dimensions of input features for all models are 3328.
@@
Due to space limitations, we present the results and analysis for MLP models in $Scene_{1}$ as a representative example in the main text. Figure 3 illustrates the fitted BNSL curve, along with the original data points. The x-axis represents the scaling factor (FLOPs), and the y-axis represents the dependent variable ($R/R^{*}$). Each data point corresponds to a model of a different size, with the number of layers and units per layer annotated in the figure. The $R^{2}$ value of the curve fitting is 0.996, indicating a strong correlation between the predicted values and the observed data points. This provides compelling empirical evidence that a Broken Neural Scaling Law governs the scaling behavior of MLP models in the typical online advertisement retrieval scenario (see Section 2). We observe consistent scaling behaviors across other scenarios and architectures. Such a scaling law enables us to predict model performance under different computational budgets and better understand efficiency-performance trade-offs. Comprehensive experimental setups and results for Transformer, DSSM, and MLP across both $Scene_{1}$ and $Scene_{2}$ are provided in Appendix C, along with an additional discussion on scaling laws in NLP and online advertising systems.